OpenStudy (anonymous):

solve the equation, e^y+5=9

OpenStudy (anonymous):

e^y = 4 y=ln 4

OpenStudy (anonymous):

e^y = 4 y = ln 4 = 1.386

OpenStudy (anonymous):

e^y = 4 means y = ln 4 by definition.

OpenStudy (anonymous):

i shouldve been more specific its e^(y+5)=9 does the rule still apply?

OpenStudy (anonymous):

oh = then y+5 = ln 9

OpenStudy (anonymous):

okay, good thats what i got too, i just wanted to double check my work. thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

shouldnt the answer be around 2.19? the answers im given to choose from are 7.2, .44, -2.8, and -4.05

OpenStudy (anonymous):

Set z= y+5 -> e^z = 9 -> z = ln 9 y+5 = ln 9 -> y = ln9 -5 = -2.80 to 2 dp